OpenStudy (anonymous):

-14x^3+9x^2-2 find zeros

OpenStudy (anonymous):

hi Welcome to Openstudy! For finding zeroes from a quadratic equation: \(ax^2+bx+c=0\) I suggest that you should use the quadratic formula, which is \(\LARGE x = \frac {-b \pm \sqrt {b^2 - 4ac}}{2a} \)